OpenStudy (anonymous):

loga(7/8) (a) Rewrite the logarithm as a ratio of common (log_10) logarithms

OpenStudy (anonymous):

loga(b/c)=loga(b)-loga(c)=log10(b)/log10(a)-log10(c)/log10(a). here b=7 and c=8.

OpenStudy (anonymous):

i got caught up wit hthe a...

OpenStudy (anonymous):

brackett, is tive me -0.057

OpenStudy (anonymous):

give *

OpenStudy (anonymous):

i got the result right?

OpenStudy (anonymous):

yeah...for calculatin the answer we need the value for a.

OpenStudy (anonymous):

the final result is?

OpenStudy (anonymous):

log10(7)/log10(a) -log10(8)/log10(a)

OpenStudy (anonymous):

\[log_a(\frac{7}{8})=\frac{log(\frac{7}{8})}{log(a)}\]

OpenStudy (anonymous):

yes worldboy

OpenStudy (anonymous):

loga( 7/8) = loga(7) – loga(8) =log10(7)/log10(a) -log10(8)/log10(a) in agreement with brackett. since we don't know anything about log 10(a) neither part can be simplified further.

OpenStudy (anonymous):

my apologies for a typing error earlier. I put loga(16x) when it should have been loga(7/8)